OpenStudy (anonymous):

solve this geomatric ralations probems Nestor has 260m of fencing material that he will use to enclose a rectangular lot if the lenght of the is 5m less than 3 times its width, what is the area of the lot? solve by any of these or choose : subsitution,comparison or elimination..... tnx

OpenStudy (chaise):

Perimetre = 2l + 2w 260m = 2L + 2W Where: L = 3xW - 5m 260m = 2(3xW) -5 +2W 260m = 8w -5 265 = 8w w = 33.125 if: 260m = 2L + 2W 260m = 2L + 2x33.125 193.75 = 2L L = 96.875 A = L x W A = 96.875 x 33.125 The area is equal to approx 3209m^2. I hope this is right. :)

OpenStudy (anonymous):

is this substitution?

OpenStudy (chaise):

This is substitution, do you follow how I got the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

how to check this?